Testing on April 15, 2014

Goals

  • Test new BOF file has fixed mode 1 ripple problem
  • Test Jim Braatz's scripts

Project info

  • TGBT13B _502_52 / 53

TGBT13B_502_52 Log

  • Scan number was not reset to 1...
  • 3pm Joe edits config file to use new BOF file, and restarts manager
  • Jason cables up the noise source to the ADC inputs.
  • Input level is around -13dB for each ADC
  • 3:30pm RippleQbandmode1 - scan 68 - ten minute integration
  • 3:40pm RippleQbandmode2 - scan 69 - ten minute integration
  • 3:50pm RippleQbandmode3 - scan 70 - ten minute integration

TGBT13B_502_53 Log

  • Scan number reset to 1
  • Fails because of leftover vegas.vfreq values lodged in Astrid (even though I exited Astrid). Then I run out of time...
Topic revision: r2 - 2015-01-27, RichardPrestage
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback